the product of two numbers is 32 and their uotient is 8. find the numbers.
Answer by ptaylor(2198)

Let x=one number
And y=the other number
We are told the following:
xy=32-------------------eq1
x/y=8-----------------eq2
From eq1, we see that x=8y. Substitute this into eq1:
8y*y=32 or
8y^2=32 divide each side vy 8
y^2=4 take square root of both sides
y=+or-2
substitute y=+or-2 into eq1:
x=+or-16 when y=+or-2
CK
(-16)*(-2)=32
32=32
and
(-16)/(-2)=8
8=8
By the same token, +16 & +2 also works
Hope this helps---ptaylor